Find the slope and the y-intercept of the line.
6x=5y+30
The slope is_______. (Type an integer or a fraction.)
The y-intercept is (0,____). (Type an integer or a fraction.)

Answer by nerdybill(7384)

Find the slope and the y-intercept of the line.
6x=5y+30
Solve by rearranging the above into "slope-intercept" form of
y = mx + b
.
6x=5y+30
6x-30=5y
(6/5)x-6=y
The slope is 6/5. (Type an integer or a fraction.)
The y-intercept is (0,-6). (Type an integer or a fraction.)